OpenStudy (anonymous):

what two numbers when multiplied together equal 120, but when added together equal 26?

OpenStudy (johnweldon1993):

So lets let number 1 = x and number 2 = y Then x * y = 120 and x + y = 26 Can you solve this system of equations?

OpenStudy (johnweldon1993):

So what we're gonna do...is solve that bottom equation for 'x' and then substitute what we find into the top equation... x + y = 26 we want to solve for 'x' so subtract 'y' from both sides... x = 26 - y Since we now know that x = 26 - y...in the first equation we have x * y = 120 turns into (26 - y)(y) = 120 Distribute this out... 26y - y^2 = 120 Subtract 120 from both sides... -y^2 + 26y - 120 = 0 Quadratic...can you solve for the solutions?